OpenStudy (anonymous):

Algebra 2 help would be greatly appreciated!

OpenStudy (anonymous):

\[3^{2x}-15(3^{x})+56=0\]

OpenStudy (astrophysics):

Hey so what's it asking

OpenStudy (astrophysics):

Solve for x? :O

OpenStudy (anonymous):

Yeah solve for x

OpenStudy (madhu.mukherjee.946):

substitute 3^x as a and solve the equation as a quadratic equation

OpenStudy (madhu.mukherjee.946):

do you need more explanation ?i mean can do it on your own

OpenStudy (madhu.mukherjee.946):

@science00000 ??

OpenStudy (anonymous):

can i multiply the -15 and 3^x ???

OpenStudy (astrophysics):

No but you can simplify it -5(3^x+1)

OpenStudy (astrophysics):

\[-5(3^{x+1})\]

OpenStudy (anonymous):

don't seem to be quite understanding this :/

OpenStudy (astrophysics):

You have to know your exponent rules, 5 x 3 = 15, so we have 5(3)(3^x) you can only do this with like bases so we have \[3^{x+1}\]

OpenStudy (astrophysics):

So here is the general rule \[x^mx^n = x^{m+n}\]

OpenStudy (astrophysics):

You can factor this I suppose

OpenStudy (astrophysics):

Yes, factoring should work

OpenStudy (astrophysics):

factoring you should end up with \[(3^x-8)(3^x-7)=0\] as we are treating 3^x as a, also suggested above

OpenStudy (anonymous):

ohh this makes a little more sense now!

OpenStudy (astrophysics):

-7 x - 8 = 56 -7+-8=15

ganeshie8 (ganeshie8):

\[\large 3^{2x}-15(3^{x})+56=0\] \[\large (\color{red}{3^{x}})^2-15(\color{red}{3^{x}})+56=0\] let \(t=\color{red}{3^{x}}\) \[\large (\color{red}t{})^2-15(\color{red}{t})+56=0\] you knw how to handle a quadratic

OpenStudy (astrophysics):

-15*

OpenStudy (anonymous):

Yes!! Thanks everyone for helping! Really appreciate it.

OpenStudy (astrophysics):

You can solve for x now, it involves logarithms, you know how to do that right? \[3^x - 7 = 0\] set it to 0 and solve for x for both

OpenStudy (anonymous):

Yeah of course. Thanks again everyone!

OpenStudy (astrophysics):

Ok, yw :)